Shift Summary: 11/30/99 Tuesday Owl

Jason LaDue (jladue@cuhep.Colorado.EDU)
Tue, 30 Nov 1999 06:05:20 -0700 (MST)


Shift Summary for owl shift, 11/30/99

On Shift: M. Barrio, J. LaDue

Hours of beam available at all intensities: 5.50
Hours of beam available at nominal intensity: 5.50
Hours of E799 data (special runs not included): 5.25
Hours of special runs (laser scans, muon runs...): 0.00

Accounting of lost time:
Pipeline reboot and init.

========================================================================
Runs ended this shift:

Tapes brought to FCC this shift:

========================================================================
Accesses during shift:

========================================================================
Overview of shift:

Beam intensity is not optimal, but at least there is beam.
Several aborts and beam loss due to accelerator problems. Only thing
that we had to do was a pipeline init.
========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
15094 00:04 First beam of the run.
00:14 Beam goes away. Ch 13 says Linac Problems
00:34 Beam returns
02:03 MCR calls and says that they cannot get three
turns going, so they will be giving us as much
beam as they can manage in 2 turns.
02:40 Beam goes away to water cooling problems
03:16 More cooling problems, still no beam.
03:50 Beam returns.
03:55 Turn on Trigger Scalers after finally realizing
that they are not running.
05:03 After getting multiple "Messages from the Pipe-
line", Decide to try a pipeline init. Crate 08
fails, so reboot it and init again. Lost about
six spills.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift
DAMP tape fraction monitoring is dead. If the beam goes away for
>= 20 minutes. Stop the run and fresh start. Question, are we still sup-
posed to run reaper when this happens?

QIE comparator errors with channel 166 at frequency of 7.